You are given the following cost data:

q TFC TVC
0 25 0
1 25 7
2 25 12
3 25 18
4 25 25
5 25 34
6 25 46
7 25 62
8 25 88

a) If the price of output is $15, how many units of output will this firm produce?

b) What is the total revenue? What is the total cost?

c) Will the firm operate or shut down in the short run? in the long run? Briefly explain your answers.
Answer #1

Answer

TC=TFC+TVC

TC(0)=25+0=25

TC(1)=25+7=32 and so on

MC(n)=TC(n)-TC(n-1)
MC(n)=marginal cost of n th unit
TC(n)=Total cost of n units of output
MC(1)=32-25=7 and so on

q TFC TVC TC MC
0 25 0 25
1 25 7 32 7
2 25 12 37 5
3 25 18 43 6
4 25 25 50 7
5 25 34 59 9
6 25 46 71 12
7 25 62 87 16
8 25 88 113 26

a)

the firm produces at P=MC or the nearest lower MC

P=15 and MC=12 then Q=6 units

the firm will produce 6 units

======

b)

total revenue =P*Q=6*15=90

TC=71

c)

The total revenue is above the total cost so the firm will operate and earn an economic profit in the short run.

the economic profit will attract new firms in the market which shifts the market supply curve to the right and increases the supply up to the market price is equal to the minimum average total cost and the firms in the market earn zero economic profit.